WAEC Further Mathematics 2013 Question 33

Question 33

Probability

Two out of 10 tickets on sale for a raffle draw are winning tickets. If a guest bought two tickets, what is the probability that both tickets are winning ticket?

  1. A.1/90
  2. B.1/45
  3. C.1/20
  4. D.1/10
Show the answer and explanation

Answer: B (1/45)

Explanation

P = 2/10 × 1/9 = 1/45.